# A modern perspective on rational homotopy theory

### v1

@@ 6.1. Localization of $1$-connected spaces
    A map of $1$-connected spaces $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\geq 2$, the induced map
    $f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$
    is an isomorphism.
@@
    An $M$-localization of a $1$-connected space $Y$ is an $M$-local homotopy equivalence $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ with $M$-local target.
@@
If $f\colon Y\rightarrow Y_{\mathbb{Q}}$ is a rationalization of $Y$, then, for every $n\geq 2$, $f$ induces the isomorphism
$f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{\mathbb{Q}}\colon\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}\mathbb{Q}\xrightarrow{\cong}\pi_{n}(Y_{\mathbb{Q}})$
that discards all torsion in the $n$-th homotopy group of $Y$ in a purely topological fashion.
@@
Every $1$-connected space can be $M$-localized uniquely up to homotopy equivalence as follows.
Theorem 6.5 ([FHT01, Theorem 9.7]).
@@
    There exists a relative CW complex $\left(Y_{M^{-1}\mathbb{Z}},Y\right)$ with no 00-cells and no $1$-cells such that the inclusion $Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ is an $M$-localization of $Y$.
@@
Note that, if $Y$ is a $1$-connected CW complex, then so is $Y_{M^{-1}\mathbb{Z}}$.
6.2. A characterization of local spaces
For a multiplicative subset $M$ of $\mathbb{Z}$, we characterize $M$-local spaces in a way that supports our model-categorical arguments for the left Bousfield localization in Section 7.
For two topological spaces $X$ and $Y$, the function complex
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)$
is the simplicial set with $n$-simplices
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}}\left(X\times|\Delta[n]|,Y\right).$
In particular, observe that
$\pi_{0}\left(\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)\right)=\left[X,Y\right]$
is the set of homotopy classes of maps from $X$ to $Y$.
The following lemma is a straightforward application of the universal property of $M$-localization.
@@ 6.2. A characterization of local spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ and $Y$ be $1$-connected spaces. If $Y$ is $M$-local, then the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Y\right)$
induced by the inclusion $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
We are now ready to establish a characterization of $M$-local spaces in the language of function complexes.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
@@
If $Y$ is $M$-local, then Lemma 6.7 implies that, for $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
is a weak homotopy equivalence. Conversely, suppose that, for $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(S^{n},Y\right)$
is a weak homotopy equivalence. Let $\tau\colon S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ denote the inclusion. Then, for every map $f\colon S^{n}\rightarrow Y$, there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
@@
commutes up to homotopy. For $n\geq 2$, let $[f]$ be an element of $\pi_{n}(Y)$ represented by a map $f\colon S^{n}\rightarrow Y$, and let $m\in M$. Then, there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
[…]
commutes up to homotopy. In particular, homotopy commutativity implies that
$[f]=f_{\ast}\left([\operatorname{id}_{S^{n}}]\right)=g_{\ast}\tau_{\ast}\left([\operatorname{id}_{S^{n}}]\right)=g_{\ast}\left([\tau]\right)\in\pi_{n}(Y).$
As $S^{n}_{M^{-1}\mathbb{Z}}$ is $M$-local, there exists a unique element
$\dfrac{1}{m}\cdot[\tau]\in\pi_{n}(S^{n}_{M^{-1}\mathbb{Z}})\cong\pi_{n}(S^{n})\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\cong\mathbb{Z}\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\cong M^{-1}\mathbb{Z}.$
Therefore, we uniquely define
$\dfrac{1}{m}\cdot[f]=g_{\ast}\left(\dfrac{1}{m}\cdot[\tau]\right)\in\pi_{n}(Y).$
We conclude that the $1$-connected space $Y$ is $M$-local. ∎
As our function complexes of $1$-reduced simplicial sets are pointed, we need the pointed version of Proposition 6.8. We write $\mathcal{T}\mathrm{op}_{\ast}$ for the category of pointed spaces. For two pointed spaces $X$ and $Y$, the function complex
$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right)$
is the simplicial set with $n$-simplices
$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X\wedge|\Delta[n]|_{+},Y\right).$
Observe that
$\pi_{0}(\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Y\right))=\left[X,Y\right]_{\ast}$
is the set of pointed homotopy classes of pointed maps from $X$ to $Y$.
In terms of these function complexes, we state the pointed version of Proposition 6.8, which is proven identically.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected pointed space $Y$ is $M$-local if and only if, for every $n\geq 2$, the map
$\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(S^{n}_{M^{-1}\mathbb{Z}},Y\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(S^{n},Y\right)$
induced by the inclusion $S^{n}\rightarrow S^{n}_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
6.3. A characterization of local homotopy equivalences
For a multiplicative subset $M$ of $\mathbb{Z}$, we also characterize the $M$-local homotopy equivalences in preparation for the left Bousfield localization in Section 7.
The following lemma, in conjunction with its pointed analog, underpins our arguments.
Lemma 6.10 ([Hir03, Corollary 9.3.3(2)]).
@@ 6.3. A characterization of local homotopy equivalences
    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of simplicial sets and $Z$ is a Kan complex, then
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,Z)$
    is a weak homotopy equivalence.
@@
    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of CW complexes and $Z$ is a topological space, then
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$
    is a weak homotopy equivalence.
We provide a series of characterizations of $M$-local homotopy equivalences in terms of function complexes.
@@
    The induced map on $M$-localizations
    $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$
    is a weak homotopy equivalence.
@@
    The induced map on $M$-localizations
    $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$
    is a homotopy equivalence.
@@
    For every $M$-local space $Z$, the map
    $f_{M^{-1}\mathbb{Z}}^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)$
    is a weak homotopy equivalence.
@@
    For every $M$-local space $Z$, the map
    $f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y,Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Z\right)$
    is a weak homotopy equivalence.
@@
${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y,Z\right)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X,Z\right)}}$$\scriptstyle{f_{M^{-1}\mathbb{Z}}^{\ast}}$$\scriptstyle{f^{\ast}}$
whose vertical maps are weak homotopy equivalences by Lemma 6.7.
It remains to show that (3) and (4) are equivalent. If (3) holds, then Lemma 6.10(2) informs us that, for every ($M$-local) space $Z$, the map
$f_{M^{-1}\mathbb{Z}}^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(Y_{M^{-1}\mathbb{Z}},Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}\left(X_{M^{-1}\mathbb{Z}},Z\right)$
is a weak homotopy equivalence. Conversely, if (4) holds, then using the induced bijection of path components
$f_{M^{-1}\mathbb{Z}}^{\ast}\colon\left[Y_{M^{-1}\mathbb{Z}},Z\right]\xrightarrow{\cong}\left[X_{M^{-1}\mathbb{Z}},Z\right]$
for $Z=X_{M^{-1}\mathbb{Z}}$ and $Z=Y_{M^{-1}\mathbb{Z}}$ yields a homotopy inverse of $f_{M^{-1}\mathbb{Z}}$. ∎
It is the equivalence of (1) and (5) in Proposition 6.11 that is relevant to our work in Section 7. Again, we need the pointed version of that equivalence, which is proven identically.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A pointed map of $1$-connected CW complexes $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if and only if, for every $M$-local pointed space $Z$, the map
$f^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(Y,Z\right)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}\left(X,Z\right)$
is a weak homotopy equivalence.

### v2

@@ 6.1. Rationalization of $1$-connected spaces
    A map of $1$-connected spaces $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\geq 2$, the map $f_{\ast}\otimes\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\otimes M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\otimes M^{-1}\mathbb{Z}$ is an isomorphism.
@@
    An $M$-localization of a $1$-connected space $Y$ is an $M$-local homotopy equivalence with $M$-local target $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$.
@@
If $f\colon Y\rightarrow Y_{\mathbb{Q}}$ is a rationalization of $Y$, then, for every $n\geq 2$, $f$ induces the isomorphism $f_{\ast}\otimes\operatorname{id}_{\mathbb{Q}}\colon\pi_{n}(Y)\otimes\mathbb{Q}\xrightarrow{\cong}\pi_{n}(Y_{\mathbb{Q}})$ that discards torsion in the $n$-th homotopy group of $Y$ in a topological fashion.
The following example features often in our arguments.
@@
What is more, the rationalization of the $n$-sphere in dimensions $n\geq 2$ is a building block for the rationalization of all $1$-connected spaces; as such, it is a cornerstone of our arguments in the remainder of this paper. We do not need the details of the geometric construction at the moment, though we present them in Section 9.2.
@@
Every $1$-connected space can be $M$-localized, and thus rationalized, uniquely up to homotopy equivalence.
Theorem 6.5 ([22, Theorem 9.7]).
@@
    There exists a relative CW complex $(Y_{M^{-1}\mathbb{Z}},Y)$ with no 0-cells and no $1$-cells such that the inclusion $Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ is an $M$-localization of $Y$.
@@
For future reference, note that, if $Y$ is a $1$-connected CW complex, then so is $Y_{M^{-1}\mathbb{Z}}$.
6.2. A characterization of rational spaces
For a multiplicative subset $M$ of $\mathbb{Z}$, we characterize $M$-local spaces, and thus rational spaces, in a way that supports our left Bousfield localization in Section 7. Namely, we describe when a $1$-connected space is $M$-local in the language of function complexes.
For two topological spaces $X$ and $Y$, the function complex $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Y)$ is the simplicial set with $n$-simplices $\operatorname{Hom}_{\mathcal{T}\mathrm{op}}(X\times|\Delta[n]|,Y)$. Observe that
$\pi_{0}(\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Y))=[X,Y]$
is the set of homotopy classes of maps from $X$ to $Y$. The following lemma is a straightforward application of the universal property of $M$-localization.
@@ 6.2. A characterization of rational spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $X$ and $Y$ be $1$-connected spaces. If $Y$ is $M$-local, then the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Y)$ is a weak equivalence.
We are now ready to characterize $M$-local spaces, and thus rational spaces, using function complexes.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence.
@@
If $Y$ is $M$-local, then Lemma 6.7 implies that the restriction map is a weak equivalence. Conversely, suppose that, for $n\geq 2$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence. For $n\geq 2$, let $[f]$ be an element of $\pi_{n}(Y)$ represented by a map $f\colon S^{n}\rightarrow Y$, and let $m\in M$. Then, there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
@@
commutes up to homotopy, where $\tau$ denotes the inclusion. In particular, homotopy commutativity implies that
$f]=f_{\ast}([\operatorname{id}_{S^{n}}])=g_{\ast}\tau_{\ast}([\operatorname{id}_{S^{n}}])=g_{\ast}([\tau]).$
As $S^{n}_{M^{-1}\mathbb{Z}}$ is $M$-local, there exists a unique element $\dfrac{1}{m}\cdot[\tau]$, so we uniquely define $\dfrac{1}{m}\cdot[f]=g_{\ast}\left(\dfrac{1}{m}\cdot[\tau]\right)$. ∎
As our function complexes of $1$-reduced simplicial sets are pointed, we need the pointed version of Proposition 6.8. We write $\mathcal{T}\mathrm{op}_{\ast}$ for the category of pointed spaces. For two pointed spaces $X$ and $Y$, the function complex $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Y)$ is the simplicial set with $n$-simplices $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Y)_{n}=\operatorname{Hom}_{\mathcal{T}\mathrm{op}_{\ast}}(X\wedge|\Delta[n]|_{+},Y)$. Again, $\pi_{0}(\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Y))=[X,Y]_{\ast}$ is the set of pointed homotopy classes of pointed maps from $X$ to $Y$. We state the pointed version of Proposition 6.8, which is proven identically.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A $1$-connected pointed space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(S^{n},Y)$ is a weak equivalence.
6.3. A characterization of rational homotopy equivalences
For a multiplicative subset $M$ of $\mathbb{Z}$, we also characterize the $M$-local homotopy equivalences using function complexes in preparation for the left Bousfield localization in Section 7. The following lemma, together with its pointed analog, underpins our arguments.
Lemma 6.10 ([26, Corollary 9.3.3(2)]).
@@ 6.3. A characterization of rational homotopy equivalences
    If $f\colon X\rightarrow Y$ is a weak equivalence of simplicial sets and $Z$ is a Kan complex, then $\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(f,Z)\colon\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{SS}\mathrm{ets}}(X,Z)$ is a weak equivalence.
@@
    If $f\colon X\rightarrow Y$ is a weak homotopy equivalence of CW complexes and $Z$ is a topological space, then $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f,Z)\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$ is a weak equivalence.
We characterize the $M$-local homotopy equivalences, and thus the rational homotopy equivalences, as follows.
@@
    The induced map on $M$-localizations $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$ is a weak homotopy equivalence.
@@
    The induced map on $M$-localizations $f_{M^{-1}\mathbb{Z}}\colon X_{M^{-1}\mathbb{Z}}\rightarrow Y_{M^{-1}\mathbb{Z}}$ is a homotopy equivalence.
@@
    For every $M$-local space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f_{M^{-1}\mathbb{Z}},Z)$ is a weak equivalence.
@@
    For every $M$-local space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f,Z)$ is a weak equivalence.
@@
$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y_{M^{-1}\mathbb{Z}},Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X_{M^{-1}\mathbb{Z}},Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(Y,Z)$$\operatorname{Map}_{\mathcal{T}\mathrm{op}}(X,Z)$}}
whose vertical maps are weak equivalences by Lemma 6.7. It remains to show that (3) and (4) are equivalent. If (3) holds, then Lemma 6.10(2) implies that, for every space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(f_{M^{-1}\mathbb{Z}},Z)$ is a weak equivalence. Conversely, if (4) holds, then using the bijection of path components $f_{M^{-1}\mathbb{Z}},Z]\colon[Y_{M^{-1}\mathbb{Z}},Z]\xrightarrow{\cong}[X_{M^{-1}\mathbb{Z}},Z]$ for $Z=X_{M^{-1}\mathbb{Z}}$ and $Z=Y_{M^{-1}\mathbb{Z}}$ yields a homotopy inverse of $f_{M^{-1}\mathbb{Z}}$. ∎
It is the equivalence of (1) and (5) that is relevant to our work in Section 7, namely its pointed version.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A pointed map of $1$-connected CW complexes $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if and only if, for every $M$-local pointed space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(f,Z)\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(Y,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(X,Z)$ is a weak equivalence.